# Dynamical systems defined by polynomials with algebraic properties

### v4

@@ 5 Stream zeros and the automorphism group
$M_{P}:=\left(\begin{array}[]{ccccc}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \cdots&\cdots&\cdots&\ddots&\cdots\\ 0&0&0&\cdots&1\\ -a_{k}&-a_{k-1}&-a_{k-2}&\cdots&-a_{1}\end{array}\right).$
Note that $M_{P}$ is not the companion matrix of the polynomial $P(z)$. Indeed, it is the companion matrix of the reciprocal polynomial of $P(z)$.
@@
The set of $k\times k$ integer matrices B such that $BM_{P}=M_{P}B$ and $\det B=\pm 1$ is denoted by $Saut_{P}$.
Recall that a square $k\times k$ integer matrix having determinant ±$\pm 1$ is called unimodular matrix. Equivalently, it is an integer matrix that is invertible over the integers. In group theory, the $k\times k$ unimodular matrices under matrix multiplication form a group called the $k\times k$ general linear group over $\mathbb{Z}$, which is denoted $GL(k,\mathbb{Z})$. Therefore,
$Saut_{P}=\{B\in GL(k,\mathbb{Z}):BM_{P}=M_{P}B\}.$
Observe that $Saut_{P}$ is a subgroup of $GL(k,\mathbb{Z})$. In fact, when $BM_{P}=M_{P}B$, then
$B^{-1}M_{P}=B^{-1}M_{P}BB^{-1}=B^{-1}BM_{P}B^{-1}=M_{P}B^{-1}.$
Furthermore, if $B_{1}M_{P}=M_{P}B_{1}$ and $B_{2}M_{P}=M_{P}B_{2}$, then
$B_{1}B_{2}M_{P}=B_{1}M_{P}B_{2}=M_{p}B_{1}B_{2}.$
Note that the $k\times k$ identity matrix $I_{k}$ always be in $Saut_{P}$, while the matrix $M_{p}$ may not be in $GL(k,\mathbb{Z})$, if $M_{p}\in GL(k,\mathbb{Z})$, then $Saut_{P}$ contains all powers of $M_{p}$ and their reflections. This happens to constitute all the elements of $Saut_{P}$ in some special cases, see the Example 4 below.
@@
Assume $(3)$. Then $\varphi:\Omega_{P}\to\Omega_{P}$ is a strong automorphism if and only if there exists a $k\times k$ integer matrix $B\in Saut_{P}$ such that
$\left(\begin{array}[]{c}y_{n+1}\\ y_{n+2}\\ \cdot\\ \cdot\\ y_{n+k}\end{array}\right)=B\left(\begin{array}[]{c}x_{n+1}\\ x_{n+2}\\ \cdot\\ \cdot\\ x_{n+k}\end{array}\right)\mbox{for any $n\in\mathbb{Z}$,}$ (4)
and $(x_{n};n\in\mathbb{Z}),(y_{n};n\in\mathbb{Z})\in\Omega_{P}$ with $\varphi(x_{n};n\in\mathbb{Z})=(y_{n};n\in\mathbb{Z})$.
@@
Let $B\in Saut_{P}$ and satisfying (4). By the hypothesis of $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$, we have
$\left(\begin{array}[]{c}y_{n+1}\\ y_{n+2}\\ \cdot\\ \cdot\\ y_{n+k}\end{array}\right)=B\left(\begin{array}[]{c}x_{n+1}\\ x_{n+2}\\ \cdot\\ \cdot\\ x_{n+k}\end{array}\right)=BM_{P}\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\\ \cdot\\ \cdot\\ x_{n+k-1}\end{array}\right)\\ =M_{P}B\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\\ \cdot\\ \cdot\\ x_{n+k-1}\end{array}\right)=M_{P}\left(\begin{array}[]{c}y_{n}\\ y_{n+1}\\ \cdot\\ \cdot\\ y_{n+k-1}\end{array}\right),$
thus $(y_{n};n\in\mathbb{Z})\in\Omega_{P}$ is well defined. On the other hand, we already know that $B^{-1}$ exists and $B^{-1}M_{P}=M_{P}B^{-1}.$ So this correspondence is clearly a strong automorphism which will be denoted as $\varphi_{B}$.
Conversely, if $\varphi$ is a strong automorphism and $(y_{n};n\in\mathbb{Z})=\varphi(x_{n};n\in\mathbb{Z})$, then $(x_{i};n\leq i\leq n+k-1)$ determines $(x_{i};i\geq n)$ which determines $(y_{i};i\geq n)$, and hence $(y_{i};n\leq i\leq n+k-1)$. This converse also holds, and the correspondence between $(x_{i};n\leq i\leq n+k-1)\in\mathbb{T}^{k}$ and $(y_{i};n\leq i\leq n+k-1)\in\mathbb{T}^{k}$ is automorphism, it is described as $\varphi_{B}$ with $B\in Saut_{P}$. ∎
Due to Theorem 4, we turn to investigate the structure of $Saut_{P}$.
@@
Assume that all roots $\theta_{1},\cdots,\theta_{k}$ of $P(z)=a_{k}z^{k}+a_{k-1}z^{k-1}+\cdots+a_{1}z+1\in\mathbb{Z}[z],k\geq 1,a_{k}\neq 0$ are simple, then we have the following:
(i) For a $k\times k$ matrix $B$, $B\in Saut_{P}$ if and only if there exist $c_{i}\in\mathbb{Z}~{}(i=0,1,\cdots,k-1)$ such that for $\lambda_{i}=c_{0}+c_{1}\theta_{i}^{-1}+\cdots+c_{k-1}\theta_{i}^{-(k-1)}~{}(i=1,\cdots,k)$ with $\lambda_{1}\lambda_{2}\cdots\lambda_{k}=\pm 1$,
$B=\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)\left(\begin{array}[]{ccc}\lambda_{1}&0&\cdots\\ 0&\ddots&0\\ \cdots&0&\lambda_{k}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)^{-1}$ (5)
holds. Hence, $Saut_{P}$ is an abelian group.
(ii) Assume additionally that all the roots of $P(z)=0$ are real, then for any $B\in Saut_{P}$, the order of $B$ is either $1,2$ or infinity. The set of $B\in Saut_{P}$ with finite order is a finite subgroup ${\mathcal{F}}_{P}$ of $Saut_{P}$, and the quotient group $Saut_{P}/{\mathcal{F}}_{P}$ is either trivial or a countable multiplicative subroup of $\mathbb{R}_{>0}$.
@@
(i) Note that $\theta_{i}\neq 0~{}(i=1,2,\cdots,k)$ by (3). Assume $B\in Saut_{P}$. Then it is an integer matrix. Since
$M_{P}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\left(\begin{array}[]{c}\theta_{i}^{k-1}\\ \vdots\\ \theta_{i}\\ 1\end{array}\right)=\theta_{i}^{-1}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)~{}(i=1,2,\cdots,k)$
and $BM_{P}=M_{P}B$, we have
$M_{P}B\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=BM_{P}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\theta_{i}^{-1}B\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)~{}(i=1,2,\cdots,k).$
Since each eigenspace of $M_{P}$ is 1-dimensional, there exists $\lambda_{i}\in\mathbb{C}$ such that

### v5

@@ 6 Stream zeros and the strong automorphism group
$C_{P}=\left(\begin{array}[]{ccccc}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \cdots&\cdots&\cdots&\ddots&\cdots\\ 0&0&\cdots&\cdots&1\\ -\frac{a_{k}}{a_{0}}&-\frac{a_{k-1}}{a_{0}}&\cdots&\cdots&-\frac{a_{1}}{a_{0}}\end{array}\right).$ (13)
be the companion matrix.
Since $C_{P}$ is not an integer matrix if $a_{0}\neq\pm 1$, the action of $C_{P}$ to $\mathbb{T}^{k}$ is multivalued. In fact, we generally define the action of a $(k\times k)$-matrix $F$ with rationl entries to $\mathbb{T}^{k}$ as $\pi\circ F\circ\pi^{-1}$, where $\pi:\mathbb{R}^{k}\to\mathbb{T}^{k}$ is the coordinatewise projection. We call $\pi\circ F\circ\pi^{-1}$ the induced action of $F$ and denote $\overline{F}$. For example, if $F=(p/q)I_{1}$ ($I_{k}$ being the identity matrix of degree $k$) with irreducible fraction $p/q$ and $x\in\mathbb{T}$, then
$\overline{F}x:=\pi\left(\frac{p}{q}\times\{\{x\}+i;~i\in\mathbb{Z}\}\right)=\left\{\frac{\{px\}}{q}+\frac{i}{q};~i=0,1,\cdots,q-1\right\},$
where $\{x\}$ being the representative of $x\in\mathbb{T}$ in $[0,1)$. Hence,
$\Omega_{P}=$ $\{(x_{n};n\in\mathbb{Z})\in\mathbb{T}^{\mathbb{Z}}:~a_{k}x_{n-k}+\cdots+a_{1}x_{n-1}+a_{0}x_{n}=0~(\forall n\in\mathbb{Z})\}$
$=$ $\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~\left(\begin{array}[]{c}x_{n-k}\\ \\ \cdot\\ \cdot\\ x_{n-1}\\ x_{n}\end{array}\right)\in\overline{C}_{P}\left(\begin{array}[]{c}x_{n-k-1}\\ \cdot\\ \cdot\\ x_{n-2}\\ x_{n-1}\end{array}\right)~(\forall n\in\mathcal{N})\right\}$
holds.
Lemma 5.
(i) The correspondence $F\mapsto\overline{F}$ defined on the set of $(k\times k)$-matrices with rational entries is one-to-one.
(ii) Any $(k\times k)$-integral matrix $F$ can be identified with $\overline{F}$ consistently. Moreover, for any $(k\times k)$-matrix $G$ with rational entries, we have
(1) $\overline{FG}=F\overline{G}$
(2) $\overline{GF}=\overline{G}F$.
Proof.
(i) Assume that $F\neq G$. Then, there exists $v\in\mathbb{R}^{k}$ such that (FG)v(F-G)v is a nonzero irrational vector. Let $u=\pi v$, Then, there exist $S,S^{\prime}\subset\mathbb{Q}^{k}$ such that $\overline{F}u=\pi F\pi^{-1}u=\pi(Fv+S)$ and $\overline{G}u=\pi G\pi^{-1}u=\pi(Gv+S^{\prime})$. Therefore, $\mathbb{O}\neq\pi(F-G)v\in(\overline{F}-\overline{G})u$. Hence, $\overline{F}\neq\overline{G}$.
(ii) Since $Fw$ is an integer vector if $w$ is so, for any $v\in\pi^{-1}u$ with $u\in\mathbb{T}^{k}$, $Fv\equiv F\{u\}~(\mbox{mod}~\mathbb{Z}^{k})$. That is, $\overline{F}u=\pi F\pi^{-1}u\equiv F\{u\}~(\mbox{mod}~\mathbb{Z}^{k})$. In this sense, we can identify $\overline{F}$ with $F$.
(1) Note that $\pi F\pi^{-1}\pi=\pi F$. Hence,
$\overline{FG}=\pi FG\pi^{-1}=\pi F\pi^{-1}\pi G\pi^{-1}=\overline{F}~\overline{G}=F\overline{G}.$
(2) Take an arbitrary $u\in\mathbb{T}^{k}$ and let $\pi v=u$ for some $v\in\mathbb{R}^{k}$. Since $F\mathbb{Z}^{k}=\mathbb{Z}^{k}$, it holds that
$\overline{GF}u=\pi GF\pi^{-1}u=\pi GF(v+\mathbb{Z}^{k})=\pi G(Fv+\mathbb{Z}^{k})=\pi G\pi^{-1}Fu=\overline{G}Fu.$
Thus, $\overline{GF}=\overline{G}F$ holds. ∎
Let $\Omega_{P}$ be the set of stream zero’s of $P(z)$. A mapping $\varphi:\Omega_{P}\to\Omega_{P}$ is called an automorphism of $\Omega_{P}$ if it is an isomorphism of $\Omega_{P}$ as the topological linear space over $\mathbb{Z}$ such that $\varphi\circ\sigma=\sigma\circ\varphi$, where $\sigma$ is the shift. Furthermore, we call it a strong automorphism if for any $(x_{n};n\in\mathbb{Z}),~(y_{n};n\in\mathbb{Z})\in\Omega_{P}$ with $\varphi(x_{n};n\in\mathbb{Z})=(y_{n};n\in\mathbb{Z})$,
$(y_{0},y_{1},\cdots,y_{k-1})\mbox{ is determined by }(x_{0},x_{1},\cdots,x_{k-1}),~\mbox{and}$
$(x_{0},x_{1},\cdots,x_{k-1})\mbox{ is determined by }(y_{0},y_{1},\cdots,y_{k-1}).$
Since a strong atomorphis $\varphi$ commutes with the shift, it holds for any $n\in\mathbb{Z}$ that
$(y_{n},y_{n+1},\cdots,y_{n+k-1})\mbox{ is determined by }(x_{n},x_{n+1},\cdots,x_{n+k-1}),~\mbox{and}$
$(x_{n},x_{n+1},\cdots,x_{n+k-1})\mbox{ is determined by }(y_{n},y_{n+1},\cdots,y_{n+k-1}).$
Note that the shift $\sigma$ on $\Omega_{P}$ is an automorphism, it is a strong automorphism as well only if $a_{0}=\pm 1$ and $a_{k}=\pm 1$.
@@
By $Saut_{P}$, we denote the set of $k\times k$ integer matrices B such that $BC_{P}=C_{P}B$ and $\det B=\pm 1$.
Note that $BC_{P}^{-1}=C_{P}^{-1}B$ also follows if $B\in Saut_{P}$ since
$BC_{P}^{-1}=C_{P}^{-1}C_{P}BC_{P}^{-1}=C_{P}^{-1}BC_{P}C_{P}^{-1}=C_{P}^{-1}B.$
@@
[…]
(i) For any $B\in Saut_{P}$, let $B=(b_{i,j})_{1\leq i,,j\leq k}$. Then, $b_{1,k},b_{2,k}\cdots,b_{k-1,k}$ are multiples of $a_{0}$, while $b_{k.k}$ and $a_{0}$ are coprime.
(ii) A mapping $\varphi:\Omega_{P}\to\Omega_{P}$ is a strong automorphism if and only there exists $B\in Saut_{P}$ such that for any $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $n\in\mathbb{Z}$,
$\left(\begin{array}[]{c}y_{n}\\ y_{n+1}\\ \cdot\\ \cdot\\ y_{n+k-1}\end{array}\right)=B\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\\ \cdot\\ \cdot\\ x_{n+k-1}\end{array}\right).$ (14)
holds, where $(y_{n};n\in\mathbb{Z}):=\varphi(x_{n};n\in\mathbb{Z})$. Therefore, a strong automorphism $\varphi$ can be identified with $B\in Saut_{P}$ as above.
@@
(i) Assume that $a_{0}\neq\pm 1$ since otherwise there is nothing to prove. Since $BC_{P}=C_{P}B$, by comparing their 1st raws, we have
$\frac{a_{k}}{a_{0}}b_{1,k}=b_{2,1},~b_{1,1}-\frac{a_{k-1}}{a_{0}}b_{1,k}=b_{2,2},~\cdots,~b_{1,k-1}-\frac{a_{1}}{a_{0}}b_{1,k}=b_{2,k}.$
As $B$ is an integer matrix, it follows that $\frac{a_{k}}{a_{0}}b_{1,k},\frac{a_{k-1}}{a_{0}}b_{1,k},\cdots,\frac{a_{1}}{a_{0}}b_{1,k}$ are integers. Since $P(z)$ is primitive and GCD of $a_{0},a_{1},\cdots,a_{k}$ is 1, this is possible only if $b_{1,k}$ is a multiple of $a_{0}$. In the same way, by comparing the 2nd, 3rd, … , (k1)(k-1)-th rows, it holds that $b_{1,k},b_{2,k},\cdots,b_{k-1,k}$ are multiples of $a_{0}$. Suppose that $b_{k.k}$ and $a_{0}$ are not coprime. Then, $b_{1,k},b_{2,k},\cdots,b_{k-1,k}$ and $b_{k,k}$ have common factor larger than 1. This contradicts with $\det B=\pm 1$.
(ii) Let $\varphi:\Omega_{P}\to\Omega_{P}$ be a strong automorphism. For any $X\in\Omega_{P}$ with $X=(x_{n};n\in\mathbb{Z})$, let $Y:=(y_{n};~n\in\mathbb{Z})=\varphi(x_{n};n\in\mathbb{Z})\in\Omega_{P}$. Then, for any $n\in\mathbb{Z}$,
$(y_{n},y_{n+1},\cdots,y_{n+k-1})\mbox{ is determined by }(x_{n},x_{n+1},\cdots,x_{n+k-1}),\mbox{ and}$
$(x_{n},x_{n+1},\cdots,x_{n+k-1})\mbox{ is determined by }(y_{n},y_{n+1},\cdots,y_{n+k-1}).$
Since $\varphi$ is an automorphism and $\Omega|_{\{n,n+1,\cdots,n+k-1\}}=\mathbb{T}^{k}~(\forall n\in\mathbb{Z})$, there exist integer matrices $\cdots,B_{-1},B_{0},B_{1},\cdots$ whose inverses are also integer matrices such that
$\left(\begin{array}[]{c}y_{n}\\ y_{n+1}\\ \cdot\\ \cdot\\ y_{n+k-1}\end{array}\right)=B_{n}\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\\ \cdot\\ \cdot\\ x_{n+k-1}\end{array}\right)$
holds for any $n\in\mathbb{Z}$. Since $\varphi\circ\sigma=\sigma\circ\varphi$, we have
$B_{1}\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \cdot\\ \cdot\\ x_{k}\end{array}\right)=B_{1}\left(\begin{array}[]{c}(\sigma X)_{0}\\ (\sigma X)_{1}\\ \cdot\\ \cdot\\ (\sigma X)_{k-1}\end{array}\right)=B_{1}B_{0}^{-1}\left(\begin{array}[]{c}(\sigma Y)_{0}\\ (\sigma Y)_{1}\\ \cdot\\ \cdot\\ (\sigma Y)_{k-1}\end{array}\right)\\ =B_{1}B_{0}^{-1}\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ \cdot\\ \cdot\\ y_{k}\end{array}\right)=B_{1}B_{0}^{-1}B_{1}\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \cdot\\ \cdot\\ x_{k}\end{array}\right).$
Therefore we have $B_{1}=B_{1}B_{0}^{-1}B_{1}$, and hence, $B_{1}=B_{0}$ follows. In the same way, it follows that $B_{n}=B_{0}$ for all $n\in\mathbb{Z}$. Let $B_{n}=B~(\forall n\in\mathbb{Z})$.
By (ii) of Lemma 4, it holds that
$\pi BC_{P}\pi^{-1}\left(\begin{array}[]{c}x_{0}\\ x_{1}\\ \cdot\\ \cdot\\ x_{k-1}\end{array}\right)=B\pi C_{P}\pi^{-1}\left(\begin{array}[]{c}x_{0}\\ x_{1}\\ \cdot\\ \cdot\\ x_{k-1}\end{array}\right)$
$=$ $B\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \cdot\\ \cdot\\ x_{k}+\{0,\frac{1}{a_{0}}\cdots\frac{a_{0}-1}{a_{0}}\}\end{array}\right)=\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ \cdot\\ \cdot\\ y_{k}+\{0,\frac{1}{a_{0}}\cdots\frac{a_{0}-1}{a_{0}}\}\end{array}\right)$
$=$ $\pi C_{P}\pi^{-1}\left(\begin{array}[]{c}y_{0}\\ y_{1}\\ \cdot\\ \cdot\\ y_{k-1}\end{array}\right)=\pi C_{P}B\pi^{-1}\left(\begin{array}[]{c}x_{0}\\ x_{1}\\ \cdot\\ \cdot\\ x_{k-1}\end{array}\right)$
for any $\left(\begin{array}[]{c}x_{0}\\ x_{1}\\ \cdot\\ \cdot\\ x_{k-1}\end{array}\right)\in\mathbb{T}^{k}$, where we used (i) to show the equality between the 3rd term and the 4th term.
Hence, $\overline{BC_{P}}=\overline{C_{P}B}$ holds, from which $BC_{P}=C_{P}B$ follows by (i) of Lemma 4. Thus, $B\in Saut_{P}$.
Conversely, for $B\in Saut_{P}$, let a mapping $\varphi:\Omega_{P}\to\Omega_{P}$ satisfy (14) for any $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $(y_{n};~n\in\mathbb{Z})=\varphi(x_{n};n\in\mathbb{Z})\in\Omega_{P}$. Then, $(y_{n},y_{n+1},\cdots,y_{n+k-1})$ is determined by $(x_{n},x_{n+1},\cdots,x_{n+k-1})$. Also considering $B^{-1}$, $(x_{n},x_{n+1},\cdots,x_{n+k-1})$ is determined by $(y_{n},y_{n+1},\cdots,y_{n+k-1})$. On the other hand, the relation (14) which holds independently of $n\in\mathbb{Z}$ implies $\varphi\circ\sigma=\sigma\circ\varphi$. Thus, $\varphi$ is a strong automorphism. ∎
Using Theorem 6, we turn to investigate the structure of $Saut_{P}$.
@@
[…]
(i) For a $k\times k$ matrix $B$, $B\in Saut_{P}$ if and only if there exists a unit $\lambda_{1}$ in $\mathbb{Q}(\theta_{1})$ with
$\lambda_{1}\in\bigcap_{i=1}^{k}(\mathbb{Z}\theta_{1}^{i-1}+\mathbb{Z}\theta_{1}^{i-2}+\cdots+\mathbb{Z}\theta_{1}^{i-k})$ (15)
such that
$B=\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}^{2}&\cdots&\theta_{k}^{2}\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)\left(\begin{array}[]{ccc}\lambda_{1}&\cdots&0\\ &\ddots&\\ 0&\cdots&\lambda_{k}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}^{2}&\cdots&\theta_{k}^{2}\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)^{-1}.$ (16)
[…]
(ii) $Saut_{P}$ is isomorphic to a direct product of $l$ number of cyclic groups of infinite orders with a finite group, where $l\leq k_{1}+k_{2}-1$ and $k_{1},2k_{2}$ are the numbers of real roots and complex roots of $P(z)=0$, respectively.
@@
(i) Assume $B\in Saut_{P}$. Since
$C_{P}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\left(\begin{array}[]{c}\theta_{i}^{k-1}\\ \vdots\\ \theta_{i}\\ 1\end{array}\right)=\theta_{i}^{-1}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)~(i=1,2,\cdots,k)$
and $BC_{P}=C_{P}B$, we have
$C_{P}B\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=BC_{P}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\theta_{i}^{-1}B\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)~(i=1,2,\cdots,k).$
Since the eigenspace of $C_{P}$ corresponding to each eigenvalue $\theta_{i}~(i=1,\cdots,k)$ is 1-dimensional, there exists a unique $\lambda_{i}\in\mathbb{C}$ such that